Equation of path of projectile motion
1
Expert's answer
2021-08-03T11:42:37-0400

The equation of path of projectile motion is a parabola and can be written as follows:


"y=xtan\theta-\dfrac{gx^2}{2v^2cos^2\theta},"

here, "x" is the horizontal displacement of the projectile, "y" is the vertical displacement of the projectile, "v" is the velocity of the projectile, "\theta" is the launch angle and "g" is the acceleration due to gravity.